To make low-fat cottage cheese, milk containing 4% butterfat is mixed with 12 gallons of milk containing 1% butterfat to obtain a mixture containing 2% butterfat. How many gallons of the fattier milk must be used?
Answer by TimothyLamb(4379)

x = gallons of milk containing 4% butterfat
---
0.04x + 0.01*12 = 0.02*(x + 12)
0.04x + 0.01*12 = 0.02x + 0.02*12
0.04x - 0.02x = 0.02*12 - 0.01*12
0.02x = 0.02*12 - 0.01*12
x = (0.02*12 - 0.01*12)/0.02
x = 6
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---